# Investigating Forecasting Models for Pandemic Infections Using Heterogeneous Data Sources: A 2-year Study with COVID-19

### v1

@@ Abstract
Index Terms:
epidemiology, pandemic forecasting, COVID-19, infections, machine learning.
@@ I Introduction
Emerging in December 2019, COVID-19 rapidly became one of the most disruptive health crises in modern history [3]. The virus spread rapidly across the globe, overwhelming healthcare systems and leading to an unprecedented number of infections, hospitalisations, and fatalities. Beyond health, it caused severe economic and social disruptions. Governments worldwide have implemented various non-pharmaceutical interventions, such as lockdowns, social distancing, and travel restrictions, to contain the spread. These measures, while effective in controlling transmission, have also created major challenges for economic stability and public well-being.
@@ II Related Work
Lastly, the paradigm of online incremental learning [18] has been used to capture changes in the data distribution (i.e., concept drift) in the spread of COVID-19, and adapt models in real time [19, 20]
@@ III-A Datasets
A diverse set of information from multiple datasets has also been considered, categorised into epidemiological, vaccination, policy, and weather.
@@ III-E Evaluation method and performance metrics
To evaluate and compare the proposed methods, we utilise widely accepted regression metrics, with a particular focus on relative error measurement.
@@ IV-A1 Role of infections (features)
Split MAPE (%) Features w/o Infections (Feat.)
@@ IV-A2 Role of epidemiological data
Split MAPE (%) Features w/o Epidemiological
@@ IV-A3 Role of vaccination data
Split MAPE (%) Features w/o Vaccinations
@@ IV-A4 Role of policy data
Split MAPE (%) Features w/o Stringency Index
@@ IV-A5 Role of weather data
     Split      MAPE (%)            Features            w/o Weather

### v2

@@ I Introduction
COVID-19 became one of the most disruptive health crises in modern history [3]. The virus spread rapidly across the globe, overwhelming healthcare systems and leading to an unprecedented number of infections, hospitalisations, and fatalities. Beyond health, it caused severe economic and social disruptions. Governments worldwide have implemented non-pharmaceutical interventions, such as lockdowns, and social distancing, to contain the spread. These measures, while effective in controlling transmission, have also created major challenges for economic stability and public well-being.
@@ II Related Work
Lastly, the paradigm of online incremental learning [18] has been used to capture changes in the data distribution (i.e., concept drift) in the spread of COVID-19, and adapt models in real time [19, 20].
@@ III-A Datasets
A diverse set of information from multiple datasets has also been considered, as described below.
@@ III-E Evaluation method and performance metrics
To evaluate and compare the proposed methods, we utilise widely accepted regression metrics.
@@ IV-A1 Role of infections (features)
Split MAPE (%)
Features
w/o Infections (Feat.)
@@ IV-A2 Role of epidemiological data
Split MAPE (%)
Features
w/o Epidemiological
@@ IV-A3 Role of vaccination data
Split MAPE (%)
Features
w/o Vaccinations
@@ IV-A4 Role of policy data
Split MAPE (%)
Features
w/o Stringency Index
@@ IV-A5 Role of weather data
   Split    MAPE (%)
   Features
   w/o Weather
@@ IV-B Important features for XGBoost and ARIMAX
Split MAPE (%)
Infections
Infections
(Raw+Feat.)
Important Feat.
(excl. Vacc.)